(D − 1)

2

(D

2 + 1)

2y = sin2

(

x

2

) + e

x + x


1
Expert's answer
2021-06-10T09:15:58-0400

Answer:-

"(D-1)^2(D^2+1)^2y=sin^2(x\/2)+e^x+x"


substitute "y=e^{kx}"

then

"Dy-\frac{dy}{dx}-ke^{kx}"

"D^2y=\frac{d^2y}{dx^2}=k^2e^{kx}"

"D^4y=\frac{d^4y}{dx^4}=k^4e^{kx}"

"(y''-2y'+y)(y^{4}+2y''+y)=0\\
(k^2e^{kx}-2ke^{kx}+e^{kx})(k^4e^{kx}+2k^2e^{kx}+e^{kx})=0\\
e^{kx}(k^2-2k+1)(k^4+2k^2+1)=0\\
k=1;k=i;k=-i"

So, we have three fundamental solutions:

"e^x;e^{ix};e^{-ix}"

Each of those three fundamental solutions satisfies the homogeneous equation and also any linear combination of those. Each of roots 𝑘 is double and to obtain three more fundamental solutions we need to multiply the corresponding fundamental solution by 𝑥 :

"xe^x;xsinx;xcosx\\"

So the general solution to the homogeneous equation is

"Y=Ae^x+Bxe^x+Csinx+Dxsinx+Ecosx+Fxcosx"

where 𝐴, 𝐵, 𝐶,𝐷, 𝐸, 𝐹 are arbitrary constants.

Since 

"sin^2\frac{x}{2}=\frac{1}{2}-\frac{cosx}{2}"

"(y''-2y'+y)(y^{(4)}+2y''+y)=\frac{1}{2}-\frac{cosx}{2}+e^x+x"


To get 𝑒x a particular solution

"\widetilde{y_1}=ax^2e^x"

now do second, third , fourth order differentiation of "\widetilde{y_1}" we get

"a=\frac{1}{32}\\
 \widetilde{y_1}=\frac{x^2e^x}{32}"


similarly

to get "(-\frac{cosx}{2})" a particular solution

"\widetilde{y_2}=bx^2cosx"

now do second, third , fourth order differentiation of "\widetilde{y_2}" we get

"b=0\\
 \widetilde{y_2}=0"


to get "(x+\frac{1}{2})" a particular solution

"\widetilde{y_3}=cx+d"

"\widetilde{y_3}'=c\\
 \widetilde{y_3}''= \widetilde{y_3}^{(4)}=0"

we get "(-2c+cx+d)(cx+d)=x+\frac{1}{2}"

c=0

"\widetilde{y_3}=0"


"\widetilde{y}= \widetilde{y_1}+ \widetilde{y_2}+ \widetilde{y_3}=\frac{x^2e^x}{32}"

answer

"\boxed{y=Y+ \widetilde{y}=Ae^x+Bxe^x+Csinx+Dxsinx+Ecosx+Fxcosx+\frac{x^2e^x}{32}}"